What is the power of a 29.1 MV lightning bolt that carries an electric current of 12.8 kA? PLEASE ANSWER IN W.
Answer #1

power = voltage * Current

Power = VI

in the above problem

Voltage = 29.1 MV = 29.1*10^6 V

current = 12.8 kA = 12.8*10^3 A

P = 29.1*10^6*12.8*10^3

= 3.725*10^11 W